A straight rod of length L extends from x=0 to x=L. The linear mass density of rod varies with x co-ordinate is λ=a0+b0x2. The gravitational force experienced by a point mass m at x=-a is...............


((please solve it for me))

Grade:12

2 Answers

Anil Pannikar AskiitiansExpert-IITB
13 years ago
Dear Pravesh,
F = Gm1*m2/d^2
dm1 = (ao + box^2) dx
m1 = ao * L + bo * L^3/3  this is mass of rod
m2 = m
here d = a + L/2
plug in all the values in formula.
Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.
All the best.
Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.
Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..
Askiitians Expert
Anil Pannikar
IIT Bombay
Ajudiya Janvi
7 years ago
Two metallic sphere of mass M are suspended by two strings each of length L The distance between the upper ends of strings is L. The angle which the strings will make with the vertical due to mutual attraction of sphere is (if each mass horizontally moved by a distance of L/4 due to mutual attraction)